OpenStudy (anonymous):

Jackie deposited $5 into a checking account in February. For each month following, the deposit amount was doubled. How much money was deposited in the checking account in the month of August?

OpenStudy (callisto):

amount = 5(2)^6 = 5(64) =$ 320

OpenStudy (anonymous):

wow thanks my answer was completely off! haha thanks!

OpenStudy (callisto):

welcome